OpenStudy (tomektomeknyc):

2 tan(4x) / 1 - tan (4x) tan (4x) is equal to: a) tan 2 x b) tan 4 x c) tan 8 x d) tan x

OpenStudy (raden):

you need this formula : tan(a+b) = (tan(a) + tan(b))/(1-tan(a)(tan(b))

hartnn (hartnn):

or even this would work directly :P tan 2A = 2 tan A/ (1+ tan^2 A)

OpenStudy (anonymous):

correction it is \[\tan 2A=\frac{ 2\tan A }{1-\tan ^{2}A }\]

OpenStudy (raden):

just typo from hartnn :P

OpenStudy (anonymous):

I know only typing error.

hartnn (hartnn):

thanks for correcting :)